# 0pt[Uncaptioned image] Dual loops     0pt[Uncaptioned image] Polycube     0pt[Uncaptioned image] Polycube segmentation

### v3

@@ 5 Results
We implemented the algorithm described in Section 4, which we refer to as the DualLoops algorithm, and evaluated its performance on a diverse set of input meshes. Our evaluation primarily focuses on two key aspects: (1) the direct quality by evaluating resulting polycube segmentations and (2) the indirect quality by evaluating hexahedral meshes generated using the polycube segmentations within a polycube-based hex-meshing pipeline.
Our dataset consists of a diverse set of 3D models collected from online sources, categorized by both geometric style and topological complexity. Geometrically, we distinguish between smooth and CAD models. Smooth models typically represent organic shapes, such as animals or humans, and are widely used in computer graphics applications like animation and game development. In contrast, CAD models depict mechanical components characterized by sharp edges and flat surfaces, commonly found in engineering contexts such as aerospace or automotive design. We also separate models based on genus, as our method currently requires manual initialization for inputs with genus $g>0$. To ensure a fair evaluation, results for higher-genus inputs are reported separately.
@@
We compare DualLoops with two publicly available polycube segmentation algorithms: PolyCut [LVS13] and EvoCube [DPM22]. For PolyCut, EvoCube, and RobustPolycube we used default parameters; for our method we used $N=10$, $N^{\prime}=30$, $\alpha=10$, and $\beta=0.01$.
@@ 5.1 Polycube segmentation quality
We briefly compare the resulting polycube segmentations generated by the three methods (PolyCut, EvoCube, and our DualLoops) by using the alignment criteria computed using $f_{a}(\mathcal{M})$ mentioned in Section 4. The complexity criteria (the number of loops) cannot be computed easily for the outer methods, as such, we opt for the following: we compute the the irregular polycube corners in the segmentation (the number of vertices with more than 2 different labels adjacent to it) and list the values in the column $f_{c}^{\prime}(\mathcal{M})$. Note that we cannot compute any validity criteria on the resulting polycube segmentations. The resulting polycube segmentations from our methods are guaranteed to be valid. The polycube segmentations resulting from the other methods may not be valid.
$f_{a}(\mathcal{M})$ $f_{c}^{\prime}(\mathcal{M})$
PolyCut 0,911 29
   Smooth ($g=0$) 0,843 38
   Smooth ($g>0$) 0,840 67
   CAD ($g=0$) 0,943 22
   CAD ($g>0$) 0,943 27
EvoCube 0,886 99
   Smooth ($g=0$) 0,797 296
   Smooth ($g>0$) 0,847 68
   CAD ($g=0$) 0,915 20
   CAD ($g>0$) 0,943 29
DualLoops 0,906 29
   Smooth ($g=0$) 0,836 38
   Smooth ($g>0$) 0,848 55
   CAD ($g=0$) 0,941 20
   CAD ($g>0$) 0,942 27
Table 1: Comparison of polycube segmentations from PolyCut [LVS13], EvoCube [DPM22], and our method (DualLoops).
@@ 5.2 Hexahedral meshing
To ensure a fair comparison, we evaluate the algorithms within a unified polycube-based hexahedral meshing pipeline. Specifically, we use the polycube segmentations generated by all three methods (PolyCut, EvoCube, and our DualLoops) as input for RobustPolycube [PRR22] to compute hexahedral meshes. As a result, the results for PolyCut and EvoCube may differ from those reported in prior work, where different downstream pipelines or parameter settings may be used.
The evaluation of the hexahedral meshes involves computing cell quality via the scaled Jacobian [PCS22], vertex quality via vertex irregularity [PCS22], and geometric fidelity using the Hausdorff distance [GLYL20, FXBH16, HJS14]. It is important to note that EvoCube is a randomized algorithm, as is our own. Therefore, different runs may yield different results. We report on the number of failed cases, instances where the hexahedral mesher cannot process the given segmentation (typically due to an invalid polycube segmentation). These failures are not penalized in the average quality metrics; instead, all evaluations are computed over the subset of successful outputs. For reproducibility, we briefly explain how these metrics are computed.
@@
We measure mesh irregularity (%irr) via the percentage of irregular vertices. A vertex of a hexahedral mesh is regular if it is adjacent to $2$, $4$, or $8$ hexahedral cells, and irregular otherwise. Similarity between the input model $\mathcal{M}$ and the surface of hexahedral mesh $\mathcal{H}$ is measured using the symmetric Hausdorff distance $\overline{\text{HD}}$. We normalize the Hausdorff distance by dividing it by the length of the diagonal of the axis-aligned bounding box encompassing $\mathcal{M}$ (the scale of $\mathcal{M}$). We further multiply the Hausdorff distance by $100$ to facilitate a concise presentation of the results.
The hexahedral meshes are visualized using HexaLab[BTP19], with the hex quality color setting showing the Scaled Jacobian of the cells.
5.3 Discussion
Quantitative results can be found in Table 1 (quality of polycube segmentations) and Table 2 (quality of resulting hexahedral meshes). The resulting polycube segmentations from PolyCut and our method (DualLoops) are of mostly similar quality. The polycube segmentations by EvoCube have a slightly lower alignment, and a higher complexity. Recall that our method is guaranteed to always return a valid polycube segmentation. This results in the hex-meshing pipeline to always succeed when given our polycube segmentations. The hex-meshing pipeline failed for some results of PolyCut (7) and EvoCube (3). Furthermore, the hexahedral meshes resulting from our polycube segmentations often are of higher quality than those constructed via the other methods in terms of the Scaled Jacobian, irregularity, and similarity (measured through the Hausdorff distance).
Qualitative results can be seen in Figure 7 (comparison of polycube segmentations and resulting hexahedral meshes) and Figure 8 (showcase of dual loops, polycube segmentations and polycubes). For models that contain features well-aligned with the principal axes, all three methods result in similar polycube segmentations, and as a result, similar hexahedral meshes. On the contrary, our algorithm sometimes fails to capture features well that are not aligned with any of the principal directions, such as the horns of goathead. Small features, such as the hands and toes of armdillo and homer, require multiple specific loops to be captured well.
[!] $\text{SJ}_{\text{min}}$ $\text{SJ}_{\text{avg}}$irr $\overline{\text{HD}}$
@@ 5.3 Discussion
Table 2: Comparison of hexahedral meshes constructed from polycube segmentations from PolyCut [LVS13], EvoCube [DPM22], and our method (DualLoops), via the polycube-based hexahedral mesh pipeline RobustPolycube [PRR22].
6 Conclusion
We presented an iterative algorithm for constructing polycube segmentations of 3D models, guaranteed to produce valid solutions, even for inputs with higher genus. Our method leverages a recent dual characterization of polycubes [SSV25] to initialize a simple, valid polycube of the appropriate genus and iteratively refine it alongside its dual loop structure. Throughout the refinement process, the validity of the solution is maintained. Our experiments demonstrate that the final polycube segmentations match or exceed the quality of state-of-the-art approaches.
A key limitation of our method is the need for a manual initialization step on high-genus inputs. While we believe that an automated initialization procedure is feasible, we do not yet have a robust solution for it.
Furthermore, our work focuses solely on generating the polycube segmentation. The downstream applications may require the accompanying polycube with additional constraints, such as non-intersecting geometry, integer vertex coordinates, or fixed edge lengths. Fortunately, the dual structure naturally lends itself to polycube construction. As such, addressing these constraints and incorporating the resulting mapping distortion into our framework are natural directions for future work.
Finally, our current loop computations rely on alignment with the principal axes. While this often produces high-quality results, certain inputs may benefit from loops that slightly deviate from axis alignment. A natural extension of our method is to incorporate more expressive direction fields into the loop computation process.
@@ 6 Conclusion
$\text{SJ}_{\text{min}}=0,07$$\text{SJ}_{\text{avg}}=0,87$$\overline{\text{HD}}=2,21$
@@
$\text{SJ}_{\text{min}}=0,13$$\text{SJ}_{\text{avg}}=0,92$$\overline{\text{HD}}=1,33$
@@
$\text{SJ}_{\text{min}}=0,07$$\text{SJ}_{\text{avg}}=0,93$$\overline{\text{HD}}=1,34$
@@
$\text{SJ}_{\text{min}}=0,08$$\text{SJ}_{\text{avg}}=0,92$$\overline{\text{HD}}=1,16$
@@
$\text{SJ}_{\text{min}}=0,08$$\text{SJ}_{\text{avg}}=0,94$$\overline{\text{HD}}=0,80$
@@
$\text{SJ}_{\text{min}}=0,05$$\text{SJ}_{\text{avg}}=0,94$$\overline{\text{HD}}=0,72$
@@
$\text{SJ}_{\text{min}}=-0,88$$\text{SJ}_{\text{avg}}=0,87$$\overline{\text{HD}}=2,25$
@@
$\text{SJ}_{\text{min}}=0,02$$\text{SJ}_{\text{avg}}=0,84$$\overline{\text{HD}}=3,15$
@@
$\text{SJ}_{\text{min}}=0,01$$\text{SJ}_{\text{avg}}=0,92$$\overline{\text{HD}}=0,97$
@@
$\text{SJ}_{\text{min}}=0,06$$\text{SJ}_{\text{avg}}=0,91$$\overline{\text{HD}}=1,25$
@@
$\text{SJ}_{\text{min}}=0,03$$\text{SJ}_{\text{avg}}=0,95$$\overline{\text{HD}}=1,60$
@@
$\text{SJ}_{\text{min}}=0,05$$\text{SJ}_{\text{avg}}=0,94$$\overline{\text{HD}}=1,32$
@@
Figure 7: Comparison of the three methods PolyCut [LVS13], EvoCube [DPM22], and our method DualLoops. We show the input mesh, and polycube segmentation and resulting hexahedral meshes per method. We include the values for $\text{SJ}_{\text{min}}$, $\text{SJ}_{\text{avg}}$, and $\overline{\text{HD}}$ per hexahedral mesh.
@@
block
(e) *
@@
amogus
@@
(i) *
@@
homer
@@
bumpytorus
(q) *
@@
airplane2
@@
(u) *
@@
goathead
@@
B33
(ac) *
(ad) *
(ae) *
B61
(af) *
(ag) *
(ah) *
B33
(ai) *
(aj) *
Figure 8: Showcase of results generated by our DualLoops algorithm, illustrating the transformation from loop structure to polycube segmentation and corresponding polycube.

### v4

@@ 5 Results
We implemented DualCube as described in Section 4. The code is available at github.com/tue-alga/DualCube. We evaluate DualCube across three aspects: direct quality of polycube segmentations, quality of the resulting hexahedral meshes, and a stress test on challenging inputs.
Our dataset (github.com/tue-alga/meshes/tree/v1.0.1) consists of a diverse set of 3D models collected from online sources categorized by both geometric style and topological complexity. Geometrically, we distinguish between smooth and CAD models. Smooth models typically represent organic shapes, such as animals or humans, and are widely used in computer graphics applications like animation and game development. In contrast, CAD models depict mechanical components characterized by sharp edges and flat surfaces, commonly found in engineering contexts such as aerospace or automotive design. We also separate models based on genus, as our method currently requires manual initialization for inputs with genus $g>0$. For a fair overview, results for higher-genus inputs are reported separately.
@@
We compare DualCube against two publicly available polycube segmentation algorithms: PolyCut [LVS13] and EvoCube [DPM22]. Table 1 reports segmentation quality in terms of fidelity and compactness, two commonly used metrics in polycube evaluation. However, as these metrics do not always predict downstream performance, we additionally measure the quality of hexahedral meshes generated by a polycube-based hexahedral meshing pipeline (RobustPolycube [PRR22]) using as input the polycube segmentations generated by the three methods. These results are presented in Table 2. Figure 12 shows an overview of our results, and Figure 13 shows side-by-side visual comparisons of polycube segmentations and resulting hexahedral meshes generated via the three different methods. To further test robustness, we evaluate all methods on the Nightmare of Polycubes dataset (github.com/LIHPC-Computational-Geometry/nightmare_of_polycubes), which is designed to expose failure cases in polycube methods. Results on these nightmares are shown in Figure 14.
For PolyCut, EvoCube, and RobustPolycube we used default parameters; for our method we used $\alpha=10$, $\beta=0.001$, $\mu=10$, $\lambda=30$, $\tau=0.5$, $\delta_{\min}=0$, $\delta_{\max}=2$, $\delta^{\prime}_{\min}=1$, $\delta^{\prime}_{\max}=1$, and $\chi=10$. It is important to note that EvoCube is a randomized algorithm, as is our own. Therefore, different runs may yield different results.
@@ 5.1 Polycube segmentation quality
fidelity compactness
PolyCut 0,911 29,1
   Smooth ($g=0$) 0,843 37,8
   Smooth ($g>0$) 0,840 67,0
   CAD ($g=0$) 0,943 21,6
   CAD ($g>0$) 0,943 27,0
EvoCube 0,884 98,7
   Smooth ($g=0$) 0,796 296,5
   Smooth ($g>0$) 0,847 67,6
   CAD ($g=0$) 0,916 20,1
   CAD ($g>0$) 0,943 29,1
DualCube 0,907 28,6
   Smooth ($g=0$) 0,836 38,0
   Smooth ($g>0$) 0,848 55,3
   CAD ($g=0$) 0,941 20,3
   CAD ($g>0$) 0,942 27,1
Table 1: Evaluation and comparison of polycube segmentations.
We evaluate and compare the polycube segmentations of the three methods in terms of fidelity and compactness, as reported in Table 1. Fidelity is measured by the dot product between each triangle’s normal and the normal of its assigned axis-aligned polycube face, averaged over all triangle faces of the resulting polycube segmentation [DPM22]. Compactness is quantified by counting the number of corners in the segmentation [DPM22], defined as vertices incident to more than two distinct labels.
While DualCube explicitly optimizes for both fidelity and compactness, potentially biasing the comparison, both PolyCut and EvoCube also target similar objectives, either directly or through related proxy terms. We are unable to verify validity of results produced by PolyCut or EvoCube. Our method guarantees valid polycube segmentations by construction.
@@ 5.2 Hexahedral meshing
We compare hexahedral meshes generated by the polycube-based hexahedral meshing pipeline RobustPolycube [PRR22], using as input the polycube segmentations generated by the three methods: PolyCut, EvoCube, and our DualCube, see Table 2. We evaluate the meshes by measuring cell quality via scaled Jacobian [PCS22], regularity via vertex irregularity [PCS22], and similarity via the Hausdorff distance [GLYL20, FXBH16, HJS14]. Furthermore, we report on the number of failure cases, instances where the pipeline cannot process the given segmentation (typically due to an invalid polycube segmentation). These failure cases are not penalized in the average quality metrics; instead, all evaluations are computed over the subset of successful outputs. For reproducibility, we briefly explain how these metrics are computed.
@@
We measure mesh irregularity (%irr) via the percentage of irregular vertices. A vertex of a hexahedral mesh is regular if it is adjacent to $2$ or $4$ hexahedral cells (on the boundary), or $8$ hexahedral cells (in the interior), and irregular otherwise. Similarity between the input model $\mathcal{M}$ and the surface of hexahedral mesh $\mathcal{H}$ is measured using the symmetric Hausdorff distance $\overline{\text{HD}}$. We normalize the Hausdorff distance by dividing it by the length of the diagonal of the axis-aligned bounding box of $\mathcal{M}$. We multiply the Hausdorff distance by $100$ for a concise presentation of the results.
fail $\text{SJ}_{\text{min}}$ $\text{SJ}_{\text{avg}}$irr $\overline{\text{HD}}$
@@
Table 2: Evaluation and comparison of resulting hex-meshes.
5.3 Nightmare of polycubes
We evaluated all three methods also on the notoriously difficult input models from the Nightmare of Polycubes dataset; see Figure 14. The nightmares contain input models for which PolyCut and EvoCube either do not return a solution or an invalid polycube segmentation. The hexahedral meshing pipeline [PRR22] appears to handle some of these invalid polycube segmentations well. Specifically, we observe that collapsing patches (for example screw and torus_step by PolyCut and EvoCube) result in low-quality hex-meshes. However, polycubes with locally overlapping edges (for example 8connected by PolyCut and EvoCube) result in high-quality hex-meshes.
Our method expects a watertight model bounding a single volume, and thus, we cannot handle 24connected. Furthermore, for higher genus inputs we require manual initialization. For this reason, we do not handle the complex higher genus inputs: twist, knot, and helix7 (not shown). For the other models, we return strictly valid polycube segmentations. For screw, torus_step, tray_step (not shown), and helix, the quality of the resulting hexahedral meshes is higher than those based on PolyCut and EvoCube. For encrusted the quality of the hex-mesh is similar, and in addition, our polycube has no locally overlapping edges. For 7connected (not shown), 8connected, and twins (not shown), our method returns a valid polycube segmentation, however, the quality of the resulting hex-mesh is not high. Currently our method routes loops only across faces. Hence, we cannot adequately capture edges with a dihedral angle approaching 0 degrees. To remedy this shortcoming in the future, we plan to extend our method to allow loops routed directly along edges instead of only across faces.
5.4 Discussion
For most models, all three methods (PolyCut, EvoCube, and our approach, DualCube) completed within a few minutes, followed by hexahedral mesh generation via the RobustPolycube pipeline in an additional few minutes. The most time-consuming case was the rocker model, where PolyCut required 80 minutes, DualCube 25 minutes, and EvoCube 10 minutes. All experiments were performed on a standard laptop (Intel i7-11th Gen, 32 GB RAM).
Quantitative results can be found in Table 1 and 2. The resulting polycube segmentations from PolyCut and our method (DualCube) are of mostly similar quality. The polycube segmentations by EvoCube have a slightly lower fidelity, and a higher compactness. The hex-meshing pipeline failed for some results of PolyCut (7) and EvoCube (3). The hex-meshing pipeline always succeeded with our polycube segmentations, likely due to our polycube segmentations to be valid by definition. Furthermore, the hexahedral meshes resulting from our polycube segmentations often are of higher quality than those constructed via the other methods in terms of the Scaled Jacobian, irregularity, and similarity (measured through the Hausdorff distance).
Qualitative results can be seen in Figure 12 and 13. The hexahedral meshes are visualized using HexaLab[BTP19], with the hex quality color setting showing the Scaled Jacobian of the cells. For models that contain features well-aligned with the principal axes, all three methods result in similar polycube segmentations, and as a result, similar hexahedral meshes. On the contrary, our algorithm sometimes fails to capture features well that are not aligned with any of the principal directions, such as the horns of goathead. Small features, such as the hands and toes of armdillo and homer, require specific loops to be captured well. The screw_head model shows a case where the solutions produced by PolyCut and EvoCube lead to poor hexahedral meshes, most likely due to invalid topology. Our method produces a polycube segmentation with a valid corresponding polycube. While the resulting hexahedral mesh is slightly improved compared to the others, it is not optimal.
(a) *
(b) rod
(c) *
(d) *
(e) block
(f) *
(g) *
(h) amogus
(i) *
(j) *
(k) homer
(l) *
(m) *
(n) cat
(o) *
(p) *
(q) bumpytorus
(r) *
(s) *
(t) airplane2
(u) *
(v) *
(w) goathead
(x) *
(y) *
(z) B48
(aa) *
(ab) *
(ac) B33
(ad) *
(ae) *
(af) B61
(ag) *
(ah) *
(ai) bottle1
(aj) *
Figure 12: Showcase of results generated by DualCube, illustrating the transformation from loop structure to polycube segmentation and corresponding polycube.
@@ 5.4 Discussion
$\text{SJ}_{\text{min}}=\mathbf{0,07}$$\text{SJ}_{\text{avg}}=\mathbf{0,87}$$\overline{\text{HD}}=\mathbf{2,21}$
@@
$\text{SJ}_{\text{min}}=\mathbf{0,13}$$\text{SJ}_{\text{avg}}=0,92$$\overline{\text{HD}}=1,33$
@@
$\text{SJ}_{\text{min}}=0,07$$\text{SJ}_{\text{avg}}=\mathbf{0,93}$$\overline{\text{HD}}=1,34$
@@
$\text{SJ}_{\text{min}}=0,08$$\text{SJ}_{\text{avg}}=0,92$$\overline{\text{HD}}=\mathbf{1,16}$
@@
$\text{SJ}_{\text{min}}=\mathbf{0,08}$$\text{SJ}_{\text{avg}}=\mathbf{0,94}$$\overline{\text{HD}}=0,80$
@@
$\text{SJ}_{\text{min}}=0,05$$\text{SJ}_{\text{avg}}=\mathbf{0,94}$$\overline{\text{HD}}=\mathbf{0,72}$
@@
$\text{SJ}_{\text{min}}=-0,88$$\text{SJ}_{\text{avg}}=\mathbf{0,87}$$\overline{\text{HD}}=\mathbf{2,25}$
@@
$\text{SJ}_{\text{min}}=\mathbf{0,02}$$\text{SJ}_{\text{avg}}=0,84$$\overline{\text{HD}}=3,15$
@@
$\text{SJ}_{\text{min}}=0,01$$\text{SJ}_{\text{avg}}=\mathbf{0,92}$$\overline{\text{HD}}=\mathbf{0,97}$
@@
$\text{SJ}_{\text{min}}=\mathbf{0,06}$$\text{SJ}_{\text{avg}}=0,91$$\overline{\text{HD}}=1,25$
@@
$\text{SJ}_{\text{min}}=0,03$$\text{SJ}_{\text{avg}}=\mathbf{0,95}$$\overline{\text{HD}}=1,60$
@@
$\text{SJ}_{\text{min}}=\mathbf{0,05}$$\text{SJ}_{\text{avg}}=0,94$$\overline{\text{HD}}=\mathbf{1,32}$
@@
Figure 13: Comparison of the three methods PolyCut [LVS13], EvoCube [DPM22], and our method DualCube. We show the input mesh, and polycube segmentation and resulting hexahedral meshes per method. We include values for $\text{SJ}_{\text{min}}$, $\text{SJ}_{\text{avg}}$, and $\overline{\text{HD}}$ per hexahedral mesh.
[…]
skip=3pt
@@
(e) screw
$\text{SJ}_{\text{min}}=-1,00$$\text{SJ}_{\text{avg}}=0,91$$\overline{\text{HD}}=51,1$
@@
$\text{SJ}_{\text{min}}=-0,99$$\text{SJ}_{\text{avg}}=0,92$$\overline{\text{HD}}=55,2$
@@
$\text{SJ}_{\text{min}}=\mathbf{0,01}$$\text{SJ}_{\text{avg}}=\mathbf{0,96}$$\overline{\text{HD}}=\mathbf{20,0}$
@@
(i) torus_step
$\text{SJ}_{\text{min}}=0,06$$\text{SJ}_{\text{avg}}=\mathbf{0,96}$$\overline{\text{HD}}=38,7$
@@
$\text{SJ}_{\text{min}}=0,06$$\text{SJ}_{\text{avg}}=0,95$$\overline{\text{HD}}=62,1$
@@
$\text{SJ}_{\text{min}}=\mathbf{0,33}$$\text{SJ}_{\text{avg}}=\mathbf{0,96}$$\overline{\text{HD}}=\mathbf{25,1}$
@@
$\text{SJ}_{\text{min}}=-0,97$$\text{SJ}_{\text{avg}}=0,93$$\overline{\text{HD}}=39,1$
@@
$\text{SJ}_{\text{min}}=\mathbf{0,03}$$\text{SJ}_{\text{avg}}=\mathbf{0,94}$$\overline{\text{HD}}=\mathbf{8,99}$
@@
(q) encrusted
$\text{SJ}_{\text{min}}=0,02$$\text{SJ}_{\text{avg}}=\mathbf{0,98}$$\overline{\text{HD}}=17,0$
@@
$\text{SJ}_{\text{min}}=-0,68$$\text{SJ}_{\text{avg}}=0,97$$\overline{\text{HD}}=32,1$
@@
$\text{SJ}_{\text{min}}=\mathbf{0,10}$$\text{SJ}_{\text{avg}}=\mathbf{0,98}$$\overline{\text{HD}}=\mathbf{16,8}$
@@
(u) 8connected
$\text{SJ}_{\text{min}}=\mathbf{0,04}$$\text{SJ}_{\text{avg}}=\mathbf{0,98}$$\overline{\text{HD}}=\mathbf{14,7}$
@@
$\text{SJ}_{\text{min}}=-0,92$$\text{SJ}_{\text{avg}}=0,97$$\overline{\text{HD}}=53,5$
@@
$\text{SJ}_{\text{min}}=\mathbf{0,04}$$\text{SJ}_{\text{avg}}=0,90$$\overline{\text{HD}}=44,0$
@@
Figure 14: Comparison of the three methods PolyCut [LVS13], EvoCube [DPM22], and our method DualCube on the nightmares. We show the input mesh, and polycube segmentation and resulting hexahedral meshes per method. We include the values for $\text{SJ}_{\text{min}}$, $\text{SJ}_{\text{avg}}$, and $\overline{\text{HD}}$ per hexahedral mesh.
6 Conclusion
We presented DualCube: an iterative algorithm for constructing polycube segmentations of 3D models. The algorithm is guaranteed to produce valid solutions, even for inputs with higher genus. Our method leverages a recent dual characterization of polycubes [SSV25] to initialize a simple, valid polycube of the appropriate genus and iteratively refine it alongside its dual loop structure. Throughout the refinement process, the validity of the solution is maintained. Our experiments demonstrate that the final polycube segmentations match or exceed the quality of state-of-the-art approaches.
A key limitation of our method is the need for a manual initialization on high-genus inputs. While we believe that a fully automated initialization is possible, but we do not yet have a solution.
Our current loop computations are based on alignment with the principal axes, which generally yields high-quality results. However, this approach has two notable limitations. First, the loop computation (and resulting polycube segmentation) is sensitive to the input mesh’s orientation. Although this sensitivity could be mitigated by re-orienting the mesh using methods such as Principal Component Analysis (PCA), some downstream applications require a fixed global orientation, making such re-orientation impractical. Second, some input meshes may benefit from loops that deviate slightly from strict axis alignment. For example, features that are not well aligned with any principal axis, even after re-orientation, may be more effectively captured using loops that adapt to local rather than global directions. A natural extension of our method is to incorporate more expressive direction fields into the loop computation process.
Lastly, our work focuses solely on generating the polycube segmentation. The downstream applications may require the accompanying polycube with additional constraints, such as non-intersecting geometry, integer vertex coordinates, or fixed edge lengths. We do not have a solution for these constraints yet, though we think the dual structure gives us the additional information needed to construct polycubes with these added constraints.